OpenStudy (anonymous):

y = 2x - 3 x + y = 3 What is the solution set? {(2, 1)} {(1, 2)} {(-1, -2)} {(-2, -1)}

OpenStudy (texaschic101):

x + y = 3 we know that y = 2x - 3....so sub 2x - 3 in for y in the equation above and solve for x. Once x is found, sub that back into either of the equations and solve for y.

OpenStudy (texaschic101):

Let me get you started... x + y = 3 x + 2x - 3 = 3 now solve for x

OpenStudy (anonymous):

x going to be 1 right

OpenStudy (texaschic101):

wrong...try again

OpenStudy (texaschic101):

x + 2x - 3 = 3 3x = 6 x = ?

OpenStudy (anonymous):

2

OpenStudy (texaschic101):

right...x = 2 now sub 2 in for x in the equation to find y x + y = 3 2 + y = 3 y = 3 - 2 y = ?

OpenStudy (anonymous):

1